# Atmospheric Density-Compensating Model Predictive Control for Targeted Reentry of Drag-Modulated Spacecraft

### v1

@@ Abstract
This paper presents an estimation and control framework that enables the targeted reentry of a drag-modulated spacecraft in the presence of atmospheric density uncertainty. In particular, an extended Kalman filter (EKF) is used to estimate the in-flight density errors relative to the atmospheric density used to generate the nominal guidance trajectory. This information is leveraged within a model predictive control (MPC) strategy to improve tracking performance, reduce control effort, and increase robustness to actuator saturation compared to the state-of-the-art approach. The estimation and control framework is tested in a Monte Carlo simulation campaign with historical space weather data. These simulation efforts demonstrate that the proposed framework is able to stay within 100 km of the guidance trajectory at all points in time for 98.4% of cases. The remaining 1.6% of cases were pushed away from the guidance by large density errors, many due to significant solar storms and flares, that could not physically be compensated for by the drag control device. For the successful cases, the proposed framework was able to guide the spacecraft to the desired location at the entry interface altitude with a mean error of 12.1 km and 99.7% of cases below 100 km.
@@ 1 Introduction
The use of small spacecraft such as CubeSats [1] has become increasingly common in recent years, enabled by the miniaturization of electronics. These small spacecraft have brought with them a decrease in the cost of building and launching a satellite into orbit. The relatively low cost these spacecraft has increased accessibility to space for educational institutions and opened up new business cases in low Earth orbit. However, due to the tight mass and volume constraints inherent in building a small spacecraft, it is difficult to incorporate a propulsion system in their design.
@@
Miniaturized propulsion systems that have been developed are expensive, which negates one of the primary benefits of building a small spacecraft - its low cost. Part of the cost is due to the fact that, to reduce the price of delivering the spacecraft to orbit, a small spacecraft typically flies as a secondary payload. To fly a propulsion system on a secondary payload, the system must undergo costly testing to ensure that it does not pose a significant risk to the primary payload. The performance of small propulsion systems is also inadequate for making significant changes to the orbit of a spacecraft [2]. As a result of these drawbacks, CubeSat designs rarely incorporate a propulsion system, resulting in a lack of maneuverability that reduces the applications for which very small spacecraft are suited.
@@
While changing the orbit of a CubeSat using propulsion may not be practical for many small satellite missions, other methods have been proposed. A spacecraft operating in low-Earth orbit (LEO) will be flying through the upper portion of Earth’s atmosphere. As the spacecraft travels through the atmosphere, a small drag force is produced, which slowly decreases the orbital energy of the spacecraft, changing the orbit until the spacecraft eventually reenters the atmosphere [3]. Researchers at NASA Ames Research Center are developing the ExoBrake, a device that enables a small spacecraft to modulate the drag force acting on the spacecraft by adjusting it’s aerodynamic properties [4]. A similar device is the Drag Deorbit Device being developed at the University of Florida [5]. Modulating the drag force allows the effect of drag to be harnessed to provide some control over the trajectory of the spacecraft without the need for propellant.
@@
The use of drag modulation by spacecraft in LEO has been investigated for several uses. One such use leverages differential drag between multiple satellites in order to fly in and maintain a formation [6, 7, 8, 9] or perform constellation phasing maneuvers [10, 11, 12]. Similarly, the rendezvous of satellites in orbit using drag modulation has also been studied [13, 14, 15, 16, 17]. Even if a spacecraft has propulsive capabilities, the use of the drag force produced during flight through an atmosphere may enable the spacecraft to change orbits using less propellant [18]. Another application of drag modulation, and the focus of this paper, is targeted reentry of spacecraft from low-Earth orbit, which can be used to safely dispose of satellites at the end of their service life and has been proposed as a method for returning small payloads to Earth [4, 19]. The HyCUBE concept [20, 21] is considering the use of drag modulation to deorbit a small, instrumented reentry vehicle in order to perform hypersonic aerothermodynamic testing.
@@
Several methods have been developed for generating nominal deorbit trajectories that target the entry interface location of a drag-modulated spacecraft [22, 23, 24, 19, 25]. These methods involve finding a sequence of ballistic coefficients as well as the times to switch between them in order to nominally reenter at a desired location relative to the surface of the Earth. However, the drag force is dependant on the density of the local atmosphere, which is highly uncertain and variable [26]. This uncertainty in the density results in uncertainty in the drag force, which would cause the spacecraft to drift far away from the nominal trajectory if the nominal control policy were implemented in an open-loop fashion. In this circumstance, it is extremely unlikely that that spacecraft would reenter near the intended location. Due to the long duration required to deorbit a spacecraft using drag and the uncertainty in the atmosphere, the spacecraft could reenter months early or late compared to the nominal trajectory.
@@
In order to compensate for atmospheric uncertainty, a new trajectory could be generated, however, these methods are computationally expensive and lack any guarantees of convergence. Additionally, if the spacecraft drifts from the nominal trajectory, a new trajectory to the desired landing site may not exist depending on how close the spacecraft is to reentry. A better option is to use the nominal trajectory as the guidance for a closed-loop control strategy. In addition to developing an algorithm for generating the nominal drag-modulated reentry trajectory, Omar [27] utilized a linear quadratic regulator to track guidance. Motivated by the expanding use of model-predictive control (MPC) for spacecraft applications [28, 29, 30], an MPC framework for drag-based deorbit tracking was developed in [31] that reduces tracking error and actuator usage by exploiting knowledge of the atmospheric density. Additionally, incremental MPC has been proposed to counter the effect of atmospheric density disturbances without the need for knowledge of the atmospheric density [32]. Previous work [31] demonstrated that MPC can provide improved tracking performance if knowledge of the atmospheric density was available but did not demonstrate a realistic method for obtaining such knowledge or provide a rigorous evaluation of the performance of the controller. And while the incremental MPC approach in [32] can counter atmospheric disturbances without knowledge of the density, the incremental formulation requires that the model be discretized over small time interval in order to preserve the validity of the assumption that the disturbance is constant over the interval. As a result, the prediction horizon of the controller is limited by computational performance, which in turn limits the performance of MPC. These limitations demonstrate the need for an MPC approach that can incorporate atmospheric density estimates derived from realistic sensor measurements and can accommodate longer timesteps than those in [32], thus enabling longer prediction horizons and improved performance.
@@
This paper presents a simple and efficient method for estimating the day-of-flight atmospheric density using a Kalman filter with GPS measurements of the motion of the spacecraft relative to the guidance trajectory. This density estimation approach is combined with an improved version of the MPC formulation in [31], which is updated to include linear time-varying (LTV) system dynamics and an MPC objective function that penalizes use of the drag device in a more practical manner. The combined estimation and control framework is exhaustively tested in this paper through a Monte Carlo simulation campaign to provide a statistical measure of the performance of the proposed framework when subjected to a wide range of initial conditions, desired reentry locations, guidance trajectories and atmospheric density dispersions that are derived from historical observations.
@@
In summary, there are four main contributions in this paper that distinguish it from the work of [31, 32]. The first is related to estimating the day-of-flight atmospheric density using a Kalman filter with GPS measurements. By estimating the atmospheric density, atmospheric disturbances can be countered without the small discretization interval required by incremental MPC [32], enabling longer prediction horizons and unlocking the inherent performance benefits of MPC. The second contribution is the extension of the MPC formulation in [31] to include LTV dynamics and a penalty on changing the drag device’s ballistic coefficient, rather than penalizing the value of the ballistic coefficient. This extension results in a more accurate prediction model and a reduction in the amount of actuation needed from the vehicle’s drag device. The third contribution is the thorough evaluation of the proposed estimation and control framework through Monte Carlo simulations using historical observations of space weather data to derive realistic atmospheric density variation. Such a simulation campaign demonstrates the effectiveness of the control framework when relying only on information that is realistically obtainable in a manner that is not considered in [31, 32]. The fourth contribution of this paper is a slightly improved method to select the nominal ballistic coefficients in the longitude targeting portion of the guidance algorithms of [24, 27]. Our proposed method helps center the guidance ballistic coefficients away from the saturation limits of the drag device, resulting in more control authority when tracking the guidance trajectory.
@@
This paper is laid out as follows: first, preliminary information regarding reference frames and notation, atmospheric density and drag, and drag modulation is provided in Section 2. Next, the methodology for generating guidance trajectories, estimating the atmospheric density and performing the model predictive control is discussed in Section 3. Subsequently, the Monte Carlo simulation campaign is described and results are presented in Section 4. Finally, the results are discussed in Section 5 followed by conclusions and future work in Section 6.
@@ 2.1.1 Earth-Centered Inertial
The Earth-centered inertial ($ECI$) frame is defined by the basis vectors $\underrightarrow{ECI}^{1}$, $\underrightarrow{ECI}^{2}$ and $\underrightarrow{ECI}^{3}$. It is accompanied by the fixed point (unforced particle), $p$, which is defined at the center of the Earth. The nonlinear dynamics of the orbiting spacecraft are expressed in the $ECI$ frame, where the position of the spacecraft is found relative to point $p$ and velocity of the spacecraft is found relative to point $p$ with time derivatives take with respect to the $ECI$ frame. Nominal spacecraft trajectories are also described in this frame. The center of mass of the spacecraft is located at a point $s$ and the position of the spacecraft relative to the center of the Earth is expressed in the $ECI$ frame as $\mbf{r}^{sp}_{ECI}$, while the velocity of the spacecraft relative to the center of the Earth with respect to the $ECI$ frame is $\mbf{v}^{sp/ECI}_{ECI}$. Guidance trajectories are similarly expressed in this frame as $\mbf{r}^{gp}_{ECI}(t)$ and $\mbf{v}^{gp/ECI}_{ECI}(t)$, where point $g$ is the center of mass of a spacecraft on the guidance trajectory.
@@ 2.1.2 Earth Centered, Earth Fixed
The Earth centered, Earth fixed ($ECEF$) frame is defined by basis vectors that rotate along with the Earth with the $\underrightarrow{ECEF}^{3}$ direction pointing North, in the direction of the rotation of the Earth. Therefore, this frame rotates at a rate of $\omega_{Earth}=7.292115\times 10^{5}$ rad/s about the $\underrightarrow{ECEF}^{3}$ direction relative to the $ECI$ frame. This frame is used to locate spacecraft relative to the surface of the Earth in order to compute the local density of the atmosphere. Additionally, it is also assumed that the atmosphere is fixed in this frame, essentially neglecting any winds. As a result, the velocity of the spacecraft expressed in the rotating $ECEF$ frame is the velocity of the spacecraft relative to the atmosphere.
@@ 2.1.3 Local Vertical, Local Horizontal
$\mbf{LVLH}^{3}_{ECI}=\frac{{\mbf{r}^{~{}gp}_{ECI}}^{\times}\mbf{v}^{~{}gp/ECI}_{ECI}}{\left\|{\mbf{r}^{~{}gp}_{ECI}}^{\times}\mbf{v}^{~{}gp/ECI}_{ECI}\right\|_{2}}.$ (2)
The $\underrightarrow{LVLH}^{2}$ direction completes a right handed coordinate system through
$\mbf{LVLH}^{2}_{ECI}={\mbf{LVLH}^{3}_{ECI}}^{\times}\mbf{LVLH}^{1}_{ECI}$ (3)
and points approximately in the in-track direction. Vectors expressed in the $ECI$ frame can be transformed to the $LVLH$ frame through multiplication by the direction cosine matrix
@@
The position of the spacecraft with center of mass at point $s$ relative to its nominal guidance position $g$, expressed in the $LVLH$ frame, is written as $\mbf{r}^{sg}_{LVLH}$ and is found through
$\mbf{r}^{sg}_{LVLH}=\mbf{r}^{sp}_{LVLH}-\mbf{r}^{gp}_{LVLH}.$ (5)
The velocity of the spacecraft relative to the guidance with respect to the $ECI$ frame and expressed in the $LVLH$ frame is written as $\mbf{v}^{sg/ECI}_{LVLH}$ and is obtained by differentiating the relative position with respect to the $LVLH$ frame using the transport theorem, giving
$\mbf{v}^{sg/ECI}_{LVLH}=\mbf{v}^{sp/LVLH}_{LVLH}-\mbf{v}^{gp/LVLH}_{LVLH}-{\boldsymbol{\omega}}_{LVLH}^{LVLH-ECI^{\times}}\mbf{r}^{sg}_{LVLH},$ (6)
@@
${\boldsymbol{\omega}}_{LVLH-ECI}=\frac{{\mbf{r}^{gp}_{LVLH}}^{\times}\mbf{v}^{gp/ECI}_{LVLH}}{{\mbf{r}^{gp}_{LVLH}}^{\mathsf{T}}\mbf{r}^{gp}_{LVLH}}$ (7)

### v2

@@ Abstract
This paper presents an estimation and control framework that enables the targeted reentry of a drag-modulated spacecraft in the presence of atmospheric density uncertainty. In particular, an extended Kalman filter is used to estimate errors between the in-flight atmospheric density density and the atmospheric density used to generate the guidance trajectory. This information is leveraged within a model predictive control strategy to improve tracking performance, reduce control effort, and increase robustness to actuator saturation compared to the state-of-the-art approach. The estimation and control framework is tested in a Monte Carlo simulation campaign with historical space weather data. These simulation efforts demonstrate that the proposed framework is able to stay within 100 km of the guidance trajectory at all points in time for 98.4% of cases. The remaining 1.6% of cases were pushed away from the guidance by large density errors, many due to significant solar storms and flares, that could not physically be compensated for by the drag control device. For the successful cases, the proposed framework was able to guide the deorbiting spacecraft to the desired location at the entry interface altitude with a mean error of 12.1 km and 99.7% of cases below 100 km.
@@ 1 Introduction
The use of small spacecraft such as CubeSats [1] has become increasingly common in recent years, enabled by the miniaturization of electronics. These small spacecraft have brought with them a decrease in the cost of building and launching a satellite into orbit. The relatively low cost of these spacecraft has increased accessibility to space for educational institutions and opened up new business cases in low Earth orbit. However, due to the tight mass and volume constraints inherent in building a small spacecraft, it is difficult to incorporate a propulsion system in their design.
@@
Miniaturized propulsion systems that have been developed are expensive, which negates one of the primary benefits of building a small spacecraft - its low cost. Small spacecraft typically fly as secondary payload to reduce delivery costs. As a result, the spacecraft must undergo extensive testing to ensure that it does not pose a significant risk to the primary payload, which drives up its overall cost. The performance of small propulsion systems is also inadequate for making significant changes to the orbit of a spacecraft [2]. For instance, as a result of these drawbacks, CubeSat designs rarely incorporate a propulsion system, resulting in a lack of maneuverability that reduces the applications for which very small spacecraft are suited.
@@
While performing orbit changes using propulsion may not be practical for many small satellite missions, other methods have been proposed. A spacecraft operating in low-Earth orbit (LEO) will be flying through the upper portion of Earth’s atmosphere in a free molecular flow regime due to the low density. As the spacecraft travels through the atmosphere, a small drag force is produced, which slowly decreases the orbital energy of the spacecraft, changing the orbit until the spacecraft eventually reenters the atmosphere [3]. Researchers at NASA Ames Research Center are developing the ExoBrake, a device that enables a small spacecraft to modulate the drag force acting on the spacecraft by adjusting it’s aerodynamic properties [4]. A similar device is the Drag Deorbit Device being developed at the University of Florida [5]. Modulating the drag force allows the effect of drag to be harnessed to provide some control over the trajectory of the spacecraft without the need for propellant.
@@
Drag modulation has been investigated to assist with maneuvering spacecraft in LEO. Examples of this include creating a differential drag between multiple satellites to fly in and maintain a formation [6, 7, 8, 9] or perform constellation phasing maneuvers [10, 11, 12]. Similarly, the rendezvous of satellites in orbit using drag modulation has also been studied [13, 14, 15, 16, 17]. Even if a spacecraft has propulsive capabilities, the use of the drag force produced during flight through an atmosphere may enable the spacecraft to use less propellant when changing orbits through an aeroassist [18] or aerocapture maneuver [19, 20]. Another application of drag modulation, and the focus of this paper, is targeted reentry of spacecraft from low-Earth orbit, which can be used to safely dispose of satellites at the end of their service life and has been proposed as a method for returning small payloads to Earth [4, 21]. The HyCUBE concept [22, 23] is considering the use of drag modulation to deorbit a small, instrumented reentry vehicle in order to perform hypersonic aerothermodynamic testing.
@@
Several methods have been developed for generating nominal deorbit trajectories that target the entry interface location of a drag-modulated spacecraft [24, 25, 26, 21, 27]. These methods involve switching between a sequence of ballistic coefficients at discrete times in order to nominally reenter at a desired location relative to the surface of the Earth. However, the drag force is dependant on the density of the local atmosphere, which is highly uncertain and variable [28]. This uncertainty in density and the corresponding drag force would cause the spacecraft to drift far away from the nominal trajectory if the nominal control policy were implemented in an open-loop fashion, likely resulting in large targeting errors at entry interface.
@@
In order to compensate for atmospheric uncertainty, a new trajectory could be generated, however, the methods in [24, 25, 26, 21, 27] are computationally expensive and lack any guarantees of convergence, making them unsuitable for real-time implementation on-board the spacecraft. Additionally, if the spacecraft drifts from the nominal trajectory, a new trajectory to the desired terminal condition (e.g., a desired latitude and longitude at entry interface) may not exist. A better option is to use the nominal trajectory as the guidance for a closed-loop control strategy. In addition to developing an algorithm for generating the nominal drag-modulated reentry trajectory, Omar [29] utilized a linear quadratic regulator to track guidance. Model predictive control (MPC) has become a practical control method for spacecraft applications [30, 31, 32], largely motivated by its ability to explicitly handle state and input constraints. A typical MPC implementation involves solving a receding-horizon optimal control problem through online optimization. Although the implementation of MPC can require substantial computational resources, careful formulation of the optimization problem (e.g., solving a nonlinear MPC problem as a linear MPC problem with approximated dynamics) and the improvement in on-board spacecraft computational capabilities have helped make MPC a practical choice for spacecraft applications [31]. This motivated the development of an MPC framework for drag-based deorbit tracking in [33] that reduced tracking error and actuator usage by exploiting knowledge of the atmospheric density. Additionally, incremental MPC has been proposed to counter the effect of atmospheric density disturbances without the need for knowledge of the atmospheric density [34]. Previous work [33] demonstrated that MPC can provide improved tracking performance if knowledge of the atmospheric density was available but did not demonstrate a realistic method for obtaining such knowledge or provide a rigorous evaluation of the performance of the controller. And while the incremental MPC approach in [34] can counter atmospheric disturbances without knowledge of the density, the incremental formulation requires that the model be discretized over small time interval in order to preserve the validity of the assumption that the disturbance is constant over the interval. As a result, the prediction horizon of the controller is limited by computational performance, which in turn limits the performance of MPC. These limitations demonstrate the need for an MPC approach that can incorporate atmospheric density estimates derived from realistic sensor measurements and can accommodate longer timesteps than those in [34], thus enabling longer prediction horizons and improved performance.
@@
This paper presents a simple and efficient method for estimating the day-of-flight atmospheric density using an extended Kalman filter (EKF) with GPS measurements of the motion of the spacecraft relative to the guidance trajectory. Inspired by atmospheric density estimation methods for aerocapture [19, 35] and atmospheric entry [36, 37], our proposed method estimates a scale factor on the nominal atmospheric density to compensate for day-of-flight density errors. This density estimation approach is combined with an improved version of the MPC formulation in [33], which is updated to include linear time-varying (LTV) system dynamics and an MPC objective function that penalizes use of the drag device in a more practical manner. The combined estimation and control framework is exhaustively tested in this paper through a Monte Carlo simulation campaign to provide a statistical measure of the performance of the proposed framework when subjected to a wide range of initial conditions, desired reentry locations, guidance trajectories and atmospheric density dispersions that are derived from historical observations.
@@
In summary, there are four main contributions in this paper that distinguish it from the work of [33, 34]. The first is related to estimating the day-of-flight atmospheric density using an EKF with GPS measurements. By estimating the atmospheric density, atmospheric disturbances can be countered without the small discretization interval required by incremental MPC [34], enabling longer prediction horizons and unlocking the inherent performance benefits of MPC. The second contribution is the extension of the MPC formulation in [33] to include LTV dynamics and a penalty on changing the drag device’s ballistic coefficient, rather than penalizing the value of the ballistic coefficient. This extension results in a more accurate prediction model and a reduction in the amount of actuation needed from the vehicle’s drag device. The third contribution is the thorough evaluation of the proposed estimation and control framework through Monte Carlo simulations using historical observations of space weather data to derive realistic atmospheric density variation. Such a simulation campaign demonstrates the effectiveness of the control framework when relying only on information that is realistically obtainable in a manner that is not considered in [33, 34]. The fourth contribution of this paper is a slightly improved method to select the nominal ballistic coefficients in the longitude targeting portion of the guidance algorithms of [26, 29]. Our proposed method helps center the guidance ballistic coefficients away from the saturation limits of the drag device, resulting in more control authority when tracking the guidance trajectory.
@@
This paper is laid out as follows: first, preliminary information regarding reference frames and notation, atmospheric density and drag, and drag modulation is provided in Section 2. Next, the methodology for generating guidance trajectories, estimating the atmospheric density and performing the model predictive control is discussed in Section 3. Subsequently, the Monte Carlo simulation campaign is described and results are presented in Section 4. Finally, conclusions and future work are given in Section 5.
@@ 2.1.1 Earth-Centered Inertial
As shown in Fig. 1, the Earth-centered inertial ($ECI$) frame is defined by the basis vectors $\underrightarrow{ECI}^{1}$, $\underrightarrow{ECI}^{2}$ and $\underrightarrow{ECI}^{3}$. It is accompanied by the fixed point (unforced particle), $p$, which is defined at the center of the Earth. The nonlinear dynamics of the orbiting spacecraft are expressed in the $ECI$ frame, where the position and velocity of the spacecraft are found relative to point $p$, with time derivatives taken with respect to the $ECI$ frame. Nominal spacecraft trajectories are also described in this frame. The center of mass of the spacecraft is located at a point $s$ and the position of the spacecraft relative to the center of the Earth is expressed in the $ECI$ frame as $\mbf{r}^{sp}_{ECI}$, while the velocity of the spacecraft relative to the center of the Earth with respect to the $ECI$ frame is $\mbf{v}^{sp/ECI}_{ECI}$. Guidance trajectories are similarly expressed in this frame as $\mbf{r}^{gp}_{ECI}(t)$ and $\mbf{v}^{gp/ECI}_{ECI}(t)$, where point $g$ is the center of mass of a spacecraft on the guidance trajectory.
@@ 2.1.2 Earth Centered, Earth Fixed
The Earth centered, Earth fixed ($ECEF$) frame is defined by basis vectors that rotate along with the Earth with the $\underrightarrow{ECEF}^{3}$ direction pointing North, in the direction of the rotation of the Earth. This frame is used to locate spacecraft relative to the surface of the Earth in order to compute the local density of the atmosphere. Additionally, it is also assumed that the atmosphere is fixed in this frame, meaning winds are neglected. As a result, the velocity of the spacecraft expressed in the rotating $ECEF$ frame is the velocity of the spacecraft relative to the atmosphere.
@@ 2.1.3 Local Vertical, Local Horizontal
$\mbf{LVLH}^{3}_{ECI}=\frac{\left({\mbf{r}^{~{}gp}_{ECI}}\right)^{\times}\mbf{v}^{~{}gp/ECI}_{ECI}}{\left\|\left({\mbf{r}^{~{}gp}_{ECI}}\right)^{\times}\mbf{v}^{~{}gp/ECI}_{ECI}\right\|_{2}},$ (2)
where $\left(\cdot\right)^{\times}$ is the skew-symmetric cross product operator [38]. The $\underrightarrow{LVLH}^{2}$ direction completes a right-handed coordinate system through $\mbf{LVLH}^{2}_{ECI}=\left({\mbf{LVLH}^{3}_{ECI}}\right)^{\times}\mbf{LVLH}^{1}_{ECI}$ and points in the in-track direction of the guidance trajectory. Vectors expressed in the $ECI$ frame can be transformed to the $LVLH$ frame through multiplication by the direction cosine matrix
@@
The position of the spacecraft with center of mass at point $s$ relative to its guidance position $g$, expressed in the $LVLH$ frame, is written as $\mbf{r}^{sg}_{LVLH}$ and is found through $\mbf{r}^{sg}_{LVLH}=\mbf{r}^{sp}_{LVLH}-\mbf{r}^{gp}_{LVLH}$. The velocity of the spacecraft relative to the guidance trajectory with respect to the $ECI$ frame and expressed in the $LVLH$ frame is written as $\mbf{v}^{sg/ECI}_{LVLH}$ and is obtained by differentiating the relative position with respect to the $LVLH$ frame using the transport theorem, giving
$\mbf{v}^{sg/ECI}_{LVLH}=\mbf{v}^{sp/LVLH}_{LVLH}-\mbf{v}^{gp/LVLH}_{LVLH}-\left({\boldsymbol{\omega}}_{LVLH}^{LVLH-ECI}\right)^{\times}\mbf{r}^{sg}_{LVLH},$ (4)
@@
${\boldsymbol{\omega}}_{LVLH-ECI}=\frac{\left({\mbf{r}^{gp}_{LVLH}}\right)^{\times}\mbf{v}^{gp/ECI}_{LVLH}}{{\mbf{r}^{gp}_{LVLH}}^{\mathsf{T}}\mbf{r}^{gp}_{LVLH}}$ (5)